OpenStudy (anonymous):

What is the value of the coefficient "a" when the quadratic equation y = (3x − 4)(2x − 1) is written in standard form? 6 5 −11 −4

OpenStudy (zzr0ck3r):

(3x-4)(2x-1) = 6x^2-3x-8x+4 = 6x^2-11x+4 ^a ^b ^c

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Is that the answer?

OpenStudy (zzr0ck3r):

what is the value of a?

OpenStudy (anonymous):

idk

OpenStudy (zzr0ck3r):

ax^2+bx+c is the form so what number is where a should be?

OpenStudy (anonymous):

uhmmm

OpenStudy (zzr0ck3r):

if the equation was 2x^2 + 3x + 7 a=2 b = 3 and c = 7 your equation is 6x^2-11x+4 so what is a?

OpenStudy (anonymous):

6

OpenStudy (zzr0ck3r):

yep